OpenStudy (fredoneedshelp):

Two times the sum of a number 9 and equals 7 .

OpenStudy (ltrout):

2x+9=7

OpenStudy (ltrout):

2(x+9)=7*

OpenStudy (fredoneedshelp):

let me check

OpenStudy (carlychrisley2202016):

well im confused about your question

OpenStudy (fredoneedshelp):

@ltrout mind helping me with a little more of these?

OpenStudy (ltrout):

Sure

OpenStudy (fredoneedshelp):

Alright ill put them here

OpenStudy (fredoneedshelp):

The difference of a number times 5 and 7 is equal to 6.

OpenStudy (ltrout):

5(x-7)=6 sounds like

OpenStudy (fredoneedshelp):

nope

OpenStudy (ltrout):

How its written is confusing. Is that exactly what it says?

OpenStudy (fredoneedshelp):

yeah.

OpenStudy (carlychrisley2202016):

hey

OpenStudy (fredoneedshelp):

yea?

OpenStudy (fredoneedshelp):

all these questions right now are translating into a multi step equation

OpenStudy (carlychrisley2202016):

this is probably a stupid question but are variables and exponents the same and the internet wont tell me.

OpenStudy (fredoneedshelp):

Exponents you use the same number like if it was 2 2 you would do 2*2

OpenStudy (carlychrisley2202016):

what???? im sorry im really stupid it confuses me.....

OpenStudy (fredoneedshelp):

Math is meant to be like that.

OpenStudy (fredoneedshelp):

even i dont understand the concept of this

OpenStudy (carlychrisley2202016):

why????

OpenStudy (fredoneedshelp):

ltrout thats how its written